The inventor of a new product believes that the cost of producing the product is given by the function C(x) = 1.75x + 7000. If the inventor charges $4 per unit, then her profit for producing and selling x units is given by the function P(x) = 2.25x – 7000
(a) What is her profit if she sells 2000 units?

(b) What is her profit if she sells 5000 units?

(c) What is the break-even point for sales stated in the number of units?

Answer by stanbon(57967)

If the inventor in the above exercise charges $4 per unit, then her profit for producing and selling units is given by the function

a. What is her profit if she sells 2000 units?
b. What is her profit if she sells 500 units
c. What is the break-even point for sales?

a. P(2000)= 2.25(2000)-7000= loss of $2500
b.P(500) = 2.25(500)-7000= loss of $5875
c.Break-even means Cost = Profit
1.75x+7000=2.25-7000
0.5x=14000
x= 28000 units